# Role of sub-barrier dynamics in coherent and incoherent radiation during tunneling ionization

### v1

@@ II Coherent radiation during tunneling ionization
We have carried out the calculation of the coherent radiation spectra in the case of a symmetric and asymmetric few-cycle laser pulses defined by their displacement
@@
where $\alpha^{\prime\prime}(t)=A^{\prime}(t)=-E(t)$, $E_{0}$ and $\omega_{0}$ are the laser field amplitude and frequency, respectively, $N=6$, and $\cal A$} is the asymmetry parameter of the laser pulse, ${\cal A}=0$ for the symmetric monochromatic pulse, and ${\cal A}=0.1$ for the asymmetric two-color one. In the calculation of the different currents $j_{ij}$, the time integrations are performed numerically, while all other integrals are evaluated analytically in the applied 3D zero-range potential case, using the corresponding ionization matrix element
$\int d^{3}\mathbf{r}\psi_{\mathbf{p}}(\mathbf{r})^{*}x\psi^{a}(\mathbf{r})=-\frac{2i\sqrt{\kappa}p_{x}}{\pi\left(\kappa^{2}+p^{2}\right)^{2}},$ (12)
where $\psi_{\mathbf{p}}$ is a plane wave with momentum $\mathbf{p}$. Further, boundary terms in the temporal Fourier transformation of the currents are treated as follows
$\int_{-\infty}^{\infty}f(s)ds=\frac{f(t_{i})^{2}}{f^{\prime}(t_{i})}+\int^{t_{f}}_{t_{i}}dsf(s)-\frac{f(t_{f})^{2}}{f^{\prime}(t_{f})}.$ (13)
@@
The spectra in a symmetric and asymmetric few-cycle laser pulses with $E_{0}=0.1$, $\omega_{0}=0.05$, and $N=6$ in the 3D zero-range potential are shown in Fig. 1. In both cases the three-step HHG dominates after third harmonic, $\omega>4$ eV, while the continuum contribution via $j_{11}$ term does at low frequencies $\omega<4$ eV, see zoom of spectra for the case of the asymmetric laser pulse in Fig. 2. The first and second harmonics are mostly due to Thomson scattering, and the SFA calculation via $j_{11}$ is well approximated by the Babushkin model in the interval $0.5<\omega<4$ eV.
In the asymmetric pulse [Fig. 2], the radiation has a spike in the near-zero-frequency domain dominated by the continuum contribution ($j_{11}$). The latter is similar to the Babushkin model. The spike at small frequencies has a simple explanation. Initially, the electron is at rest at the core, then it tunnels out during the current formation time $\tau$, and is accelerated in the continuum due to the laser field up to its final velocity $v$. The current associated with this motion can be modeled as
$j(t)\sim v\,\,[1+\text{erf}(t/\tau)],$ (14)
with the Fourier component $j_{\omega}\sim v\exp(-\omega^{2}\tau^{2}/4)/\omega$. Thus, from Eq. (5)
$\frac{dw}{d\omega}\sim\frac{v^{2}}{\omega}e^{-\omega^{2}\tau^{2}/2},$ (15)
and the spectrum has a singularity at zero-frequency.
While $j_{11}$ shows the near-zero-frequency spike in the spectrum, it however includes additionally the under-the-barrier effect, which enhances the emission in most of the spectral region (at $\gamma=0.5$ the Babushkin model exaggerates the singularity because it does not account accurately the formation time of the current. Here, the error-function is replaced by a Heaviside-theta function). The enhancement is mostly due to the acceleration of the electron during tunneling. At larger Keldysh parameters $\gamma=\kappa\omega_{0}/E_{0}$ [cf. the cases $\gamma=0.5$ and $\gamma=0.7$ in Fig. 2] the effect of the under-the barrier dynamics is stronger. In particular, the enhancement of the near-zero frequency peak with respect to the Babushkin model is larger. However, the probability of the radiation is decreased. The width of the near-zero-frequency spike is about $\Delta\omega\sim 0.2$ eV, which is related to the laser pulse duration. In fact, the FWHM of the pulse in intensity is approximately $\tau\sim 200$ a.u. which corresponds to an energy of $1/\tau\sim 0.15$ eV, roughly fitting to $\Delta\omega\sim 1/\tau$.
To demonstrate the contribution of the under-the-barrier dynamics for the coherent radiation, we present the emission probability of Eq. (5) in the following form in the 1D case:
$\frac{dw}{d\omega}=\frac{2\alpha\omega}{3\pi c^{2}}\int dx\int dyj_{\omega}(x)j^{*}_{\omega}(y),$ (16)
with the current Fourier components $j_{\omega}(x)$. The function $j_{\omega}(x)j^{*}_{\omega}(y)$ is plotted in Fig. 3. From the latter we can distinguish which coordinate regions $x\sim y$ contribute significantly in the emission probability. We see in Fig. 3 that the under-the-barrier region has nonnegligible contribution to the radiation.
@@ III Spontaneous radiation during tunneling ionization
$dw=\left|m(\mathbf{k},\mathbf{p})\right|^{2}\frac{d^{3}\mathbf{k}}{(2\pi)^{3}}\frac{d^{3}\mathbf{p}}{(2\pi)^{3}},$ (17)
@@
$m(\mathbf{k},\mathbf{p})=\int dt\langle\psi_{\mathbf{p}}^{f}(t),1|\hat{\mathbf{p}}\cdot\hat{\mathbf{A}}(t)|\Psi(t),0\rangle,$ (18)
with the SFA wave function of the tunneling electron $\Psi(t)$. Here ${\cal V}=1$ is assumed. Integrating over the final electron momenta we obtain:
$\frac{dw}{d\omega}$ $=$ $\frac{4\pi\omega^{2}}{c^{3}}\int d^{3}\mathbf{p}\left|\int dt\langle\psi_{\mathbf{p}}^{f}(t),1|\hat{\mathbf{p}}\cdot\hat{\mathbf{A}}(t)|\Psi(t),0\rangle\right|^{2}.$ (19)
In the first-order SFA $\Psi=\psi^{(0)}(t)$ and we obtain the photon emission probability which coincides with the result of Ref. [14] in velocity gauge. In the second-order SFA calculation, the electron initial state during the photon emission process is $\Psi(t)=\psi^{(0)}(t)+\psi^{(1)}(t)$, which includes the continuum dynamics in the tunneling ionization process via $\psi^{(1)}(t)$. Comparing the first-order probability with the second-order one, we learn about the role of the continuum dynamics for the spontaneous radiation. We define the zero-order amplitude as
@@
We have carried out the calculation of the incoherent radiation spectra in the case of a symmetric and asymmetric few-cycle laser pulses defined above with $E_{0}=0.1$, $\omega_{0}=0.05$, and $N=2$.
Here again, in the calculation of the amplitudes $m_{i}$ the time integrations are performed numerically. All other inner integrals are calculated analytically. Especially, $m_{1}$ has the structure of a nested integral and is transformed via two partial integrations in the following way
$m_{1}$ $=$ $\int ds\left[\mathbf{p}+\mathbf{A}(s)\right]\exp(i\omega s)\int^{s}dt\tilde{m}_{1}(t)$ (23)
$=$ $-\int ds\left[\mathbf{p}+\mathbf{A}(s)\right]\tilde{m}_{1}(s)\int^{s}dt\exp(i\omega t)$
$+\int ds\mathbf{E}(s)\int^{s}dt\tilde{m}_{1}(t)\int^{s}dt\exp(i\omega t)$
$=$ $-\int ds\left[\mathbf{p}+\mathbf{A}(s)\right]\tilde{m}_{1}(s)\int^{s}dt\exp(i\omega t)$
$-\int ds\tilde{m}_{1}(s)\int^{s}dr\mathbf{E}(r)\int^{r}dt\exp(i\omega t).$
@@
All inner time integrals are analytical and the outer numerical integration one-dimensional over the full time-space. The final momentum integration is also performed numerically. Further, boundary terms of the time integrals are treated like
$\int dsf(s)=\int^{t_{f}+2i\gamma/\omega}_{t_{i}+2i\gamma/\omega}dsf(s).$ (24)
@@
The spectra for the spontaneous radiation are shown in Fig. 4. The spectrum via $m_{0}(\mathbf{k},\mathbf{p})$ is similar to the calculation of Ref. [14], with the spectral width $\Delta\omega\sim 1/\tau_{K}$, determined by the Keldysh time $\tau_{K}=\kappa/E_{0}$. Up to $\omega<6$ eV it is much smaller with respect to $m_{1}(\mathbf{k},\mathbf{p})$, which includes the continuum motion. In the near-zero-frequency region, $m_{1}(\mathbf{k},\mathbf{p})$ has a spike similar to that in the coherent radiation. At larger frequencies $\omega>6$ eV, $m_{0}(\mathbf{k},\mathbf{p})$ and $m_{1}(\mathbf{k},\mathbf{p})$ compensate each other. Thus, the amplitude $m_{0}(\mathbf{k},\mathbf{p})$ does not describe the observable spectrum neither at low frequencies $\omega\rightarrow 0$, nor at high ones $\omega\rightarrow\infty$, but only near the cutoff $\omega\sim 8$ eV. Our remarkable observation is that there is no significant difference in spontaneous radiation in symmetric or asymmetric pulses.
The spontaneous emission spectra in Fig. 4 have been calculated integrating over the final electron momentum. However, when the correlated electron and photon measurement is carried out in an experiment, one can access the electron momentum resolved photon spectra. In Fig.5 we present the electron momentum resolved total photon yield in the case of the correlated measurement. Calculation are carried out in the case of 1D potential and a half-cycle laser pulse, using saddle-point-approximation in evaluating the emission matrix element. Small contribution of the under-the-barrier recollisions via $m_{2}$ can be distinguished near the peak. However, this does not lead to any qualitative change of the spectrum.
@@
IV Estimation of radiation intensities
Let us estimate the relation between the coherent and incoherent Brunel radiation. We use the dipole radiation power [29]:

### v2

@@ II.3.1 Calculation of the current
We have carried out the calculation of the coherent radiation spectra in the case of a symmetric and an asymmetric few-cycle laser pulse defined by their displacement
@@
where $\alpha^{\prime\prime}(t)=A^{\prime}(t)=-E(t)$, $E_{0}$ and $\omega_{0}$ are the laser field amplitude and frequency, respectively, $N=12$, $\cal A$} is the asymmetry parameter of the laser pulse, ${\cal A}=0$ for the symmetric monochromatic pulse, and ${\cal A}=0.1$ for the asymmetric two-color one. In the calculation of currents, the time integrations are performed numerically, while all other integrals are evaluated analytically in the applied 3D zero-range potential case.
The current component $\mathbf{J}_{01}(t)$ reads
$\mathbf{J}_{01}(t)=-i\int^{t}dt^{\prime}\int d^{3}\mathbf{p}\,\langle\psi^{(0)}(t)\left|\hat{\mathbf{p}}\right|\psi^{(f)}_{\mathbf{p}}(t)\rangle\langle\psi^{(f)}_{\mathbf{p}}(t^{\prime})|\mathbf{r}\cdot\mathbf{E}(t^{\prime})|\psi^{(a)}(t^{\prime})\rangle,$
which after the saddle-point integration over the momentum is expressed as follows:
$\mathbf{J}_{01}(t)=-i\int^{t}dt^{\prime}\left(\frac{2\pi i}{t-t^{\prime}-i\epsilon}\right)^{3/2}\mathbf{m}_{p}^{*}(\mathbf{p}_{s},t)m_{x}(\mathbf{p}_{s},t^{\prime})$ (16)
where $\mathbf{p}_{s}=-[\bm{\alpha}(t)-\bm{\alpha}(t^{\prime})]/(t-t^{\prime})$ is the saddle-point momentum, and the singularity is handled using $\epsilon=0.4/\kappa^{2}$ [48]. Here, the matrix elements are defined
$\mathbf{m}_{p}(\mathbf{p},t)$ $=$ $\int d^{3}\mathbf{r}\psi^{f}_{\mathbf{p}}(\mathbf{r},t)^{*}\hat{\mathbf{p}}\psi^{a}(\mathbf{r},t)=\frac{\sqrt{\kappa}(\mathbf{p}+\mathbf{A}(t))}{\pi\left[\kappa^{2}+(\mathbf{p}+\mathbf{A}(t))^{2}\right]^{2}}$
$\times\exp[ip^{2}/2t+i\mathbf{p}\cdot\bm{\alpha}(t)+i\beta(t)+i\kappa^{2}/2t],$
$m_{x}(\mathbf{p},t)$ $=$ $\int d^{3}\mathbf{r}\psi^{f}_{\mathbf{p}}(\mathbf{r},t)^{*}\hat{\mathbf{r}}\cdot\mathbf{E}(t)\psi^{a}(\mathbf{r},t)$ (17)
$=$ $-2i\,\mathbf{m}_{p}(\mathbf{p},t)\cdot\mathbf{E}(t),$
with $\beta=\int dt\,A^{2}(t)/2$.
The current component $\mathbf{J}_{11}(t)$ reads
$\mathbf{J}_{11}(t)$ $=$ $-\int^{t}dt^{\prime}\int^{t}dt^{\prime\prime}\int d^{3}\mathbf{p}\int d^{3}\mathbf{p}^{\prime}\langle\psi^{(a)}(t^{\prime})|\mathbf{r}\cdot\mathbf{E}(t^{\prime})|\psi^{(f)}_{\mathbf{p}}(t^{\prime})\rangle$ (18)
$\times$ $\langle\psi^{(f)}_{\mathbf{p}}(t)\left|\hat{\mathbf{p}}\right|\psi^{(f)}_{\mathbf{p}^{\prime}}(t)\rangle\langle\psi^{(f)}_{\mathbf{p}}(t^{\prime\prime})|\mathbf{r}\cdot\mathbf{E}(t^{\prime\prime})|\psi^{(a)}(t^{\prime\prime})\rangle,$
$=$ $\int d^{3}\mathbf{p}(\mathbf{p}+\mathbf{A}(t))\left|\int^{t}dt^{\prime}m_{x}(\mathbf{p},t^{\prime})\right|^{2}.$
The $t$-integration in Eq. (16), $t$- and $p$-integrations in Eq. (18), and the Fourier transform in Eq. (7) are calculated numerically.
Further, because of a finite duration of the laser pulse with sharp time boundaries, unphysical contributions arise [49] in the temporal Fourier transformation of the currents, which are removed by the following procedure. In the interval $(t_{i},t_{f})$ the time-integration for the Fourier transform is carried out numerically, while beyond it analytically:
$\int_{-\infty}^{\infty}dtf(t)=\frac{f(t_{i})^{2}}{f^{\prime}(t_{i})}+\int^{t_{f}}_{t_{i}}dtf(t)-\frac{f(t_{f})^{2}}{f^{\prime}(t_{f})},$ (19)
where $t_{i}$ and $t_{f}$ are the switching-on and -off times of the laser pulse, respectively, using
$\int^{t_{i}}_{-\infty}dtf(t)$ $\approx$ $\int^{t_{i}}_{-\infty}dt\exp[\ln(f(t_{i})+f^{\prime}(t_{i})(t-t_{i}))]=\frac{f(t_{i})^{2}}{f^{\prime}(t_{i})},$
with the function $f$ being the integrand of the Fourier-transformation $f(t)=J(t)\exp[i\omega t-\epsilon t]$, where the parameter $\epsilon=2.4\times 10^{-5}$ a.u. is introduced corresponding to the interaction time of 1 ps (frequency resolution of 1 THz).
Figure 3: Spontaneous radiation spectrum during tunneling ionization: (left column) asymmetric few-cycle laser pulse, (right column) symmetric laser pulse, in both case $N=4$; (a,b) for $E_{0}=0.1$, (c,d) for $E_{0}=0.075$. The contributions of amplitudes in different approximations $m_{0}$, $m_{1}$, and $m_{0}+m_{1}$ are specified. The results of the simple man Drude model are indicated by $m_{D}$.
II.3.2 Radiation spectra
@@
The spectra in symmetric and asymmetric few-cycle laser pulses with $E_{0}=0.1$, $\omega_{0}=0.05$, and the number of cycles $N=12$ in the 3D zero-range potential are shown in Fig. 1. In both cases the three-step HHG dominates after the third harmonic, $\omega>3\omega_{0}\approx 4$ eV, while the continuum contribution via $J_{11}$ term does at low frequencies $\omega<4$ eV, see zoom of spectra for the case of the asymmetric laser pulse in Fig. 2. The first and second harmonics are mostly due to Thomson scattering, and the SFA calculation via $J_{11}$ is well approximated by the Drude model in the interval $0.5<\omega<4$ eV. In particular, in the asymmetric pulse [Fig. 2], the radiation has a spike in the near-zero-frequency domain dominated by the continuum contribution ($J_{11}$), characteristic to the Drude model. Thus, we conclude that $J_{11}$ at low frequencies describes the Brunel radiation.
While Fig. 1 makes the impression that the $J_{11}$ induced radiation spectrum in the near-zero-frequency domain coincides with that via the Drude model (with the solely classical continuum contribution), the zoomed-in view of Fig. 2 demonstrates differences from the Drude model, especially at the peak of the near-zero-frequency region. The reason of the difference is that $J_{11}$ includes additionally the quantum free-free transition effect. The latter modifies the emission in most of the near-zero-frequency spectral region. The comparison of the cases $\gamma=0.5$ and $\gamma=0.7$ in Fig. 2(c,d) shows that the quantum dynamics leads to small enhancement of the THz spectrum with respect to the classical Drude model in the nonadiabatic regime and vice versa in the adiabatic regime.
The width of the near-zero-frequency spike is about $\Delta\omega\sim 0.2$ eV, which is related to the laser pulse duration. In fact, the FWHM of the pulse in intensity is approximately $\tau\sim 200$ a.u. which corresponds to an energy of $1/\tau\sim 0.15$ eV, roughly fitting to $\Delta\omega\sim 1/\tau$.
@@ III.1 Spontaneous radiation spectra
$dw=\left|m(\mathbf{k},\mathbf{p})\right|^{2}d^{3}\mathbf{k}d^{3}\mathbf{p},$ (20)
@@
$m(\mathbf{k},\mathbf{p})=\int dt\langle\psi_{\mathbf{p}}^{(f)}(t),1|\hat{\mathbf{p}}\cdot\hat{\mathbf{A}}(t)|\Psi(t),0\rangle,$ (21)
with the SFA wave function of the tunneling electron $\Psi(t)$. Integrating over the final electron momenta we obtain:
$\frac{dw}{d\omega}$ $=$ $\frac{4\pi\omega^{2}}{c^{3}}\int d^{3}\mathbf{p}\left|\int dt\langle\psi_{\mathbf{p}}^{(f)}(t),1|\hat{\mathbf{p}}\cdot\hat{\mathbf{A}}(t)|\Psi(t),0\rangle\right|^{2}.$ (22)
In zeroth-order SFA $\Psi=\psi^{(0)}(t)$ and we obtain the photon emission probability which coincides with the result of Ref. [30] in velocity gauge. In the first-order SFA calculation, the electron initial state during the photon emission process is $\Psi(t)=\psi^{(0)}(t)+\psi^{(1)}(t)$, which includes the continuum dynamics in the tunneling ionization process via $\psi^{(1)}(t)$. Comparing the zeroth-order probability with the first-order one, we learn about the role of the continuum dynamics for the spontaneous radiation. We define the zero-order amplitude as
@@
The amplitudes are expressed in the following form:
$m_{0}(\mathbf{k},\mathbf{p})$ $=$ $-i\int dt\hat{\mathbf{A}}(t)\cdot\mathbf{m}_{p}^{*}(\mathbf{p},t)$ (27)
$m_{1}(\mathbf{k},\mathbf{p})$ $=$ $\int dt^{\prime}\left[\mathbf{p}+\mathbf{A}(t^{\prime})\right]\cdot\hat{\mathbf{A}}(t^{\prime})\int^{t^{\prime}}dtm_{x}(\mathbf{p},t).$
The amplitude $m_{1}$ has a structure of a nested integral and is transformed via two partial integrations to the following expression
$m_{1}$ $=$ $-\int dt^{\prime}\left[\mathbf{p}+\mathbf{A}(t^{\prime})\right]\cdot\hat{\mathbf{A}}(t^{\prime})\,m_{x}(\mathbf{p},t^{\prime})\int^{t^{\prime}}dt\exp(i\omega t)$ (28)
$-\int dt^{\prime}m_{x}(\mathbf{p},t^{\prime\prime})\int^{t^{\prime\prime}}dt^{\prime}\mathbf{E}(t^{\prime})\int^{t^{\prime}}dt\exp(i\omega t).$
@@
The inner time integrals are analytical and the outer integration is performed numerically. The final momentum integration is also performed numerically.
The boundaries of the finite time integration introduce not physically relevant time diffraction oscillations in the integration result. To avoid it, we shift the time integration contour in the complex plane passing over the saddle point $t_{s}=i\gamma/\omega$, and omit the contribution of the vertical contours, which are responsible for the time diffraction, see a similar procedure in Ref.  [49]. Thus, the boundary terms of the time integrals are treated like
$\int dtf(t)=\int^{t_{f}+i\gamma/\omega}_{t_{i}+i\gamma/\omega}dtf(t).$ (29)
@@
We have carried out the calculation of the incoherent radiation spectra in the case of a symmetric and asymmetric few-cycle laser pulses defined above with $E_{0}=0.1$ ($\gamma=0.5$) and $E_{0}=0.075$ ($\gamma=0.7$), $\omega_{0}=0.05$, and $N=4$. The spectra for the spontaneous radiation are shown in Fig. 3. The spectrum via $m_{0}(\mathbf{k},\mathbf{p})$ is similar to the calculation of Ref. [30], with the spectral width $\Delta\omega\sim 1/\tau_{K}$, determined by the Keldysh time $\tau_{K}=\kappa/E_{0}$. Up to $\omega<6$ eV it is much smaller with respect to $m_{1}(\mathbf{k},\mathbf{p})$, which includes the continuum motion. In the near-zero-frequency region, $m_{1}(\mathbf{k},\mathbf{p})$ has a spike similar to that in the coherent radiation. At larger frequencies $\omega>6$ eV, $m_{0}(\mathbf{k},\mathbf{p})$ and $m_{1}(\mathbf{k},\mathbf{p})$ compensate each other. Thus, the amplitude $m_{0}(\mathbf{k},\mathbf{p})$ does not describe the observable spectrum neither at low frequencies $\omega\rightarrow 0$, nor at high ones $\omega\rightarrow\infty$, but only near the cutoff $\omega\sim 8$ eV. Our remarkable observation is that there is no significant difference in spontaneous radiation in symmetric or asymmetric pulses. The qualitative behavior of the spectra does not change with increasing the Keldysh parameter, but the intensity of radiation decreases with larger $\gamma$ as the driving electric field decreases. The spectral width of the radiation is related to the Keldysh time $\Delta\omega\sim 1/\tau_{K}=\gamma/\omega$.
III.2 Simple man Drude model for spontaneous radiation
For an intuitive proof of the spontaneous radiation spectra, we have generalized the simple man Drude model of Eq. (12) for the spontaneous radiation described by Eq. (22). Taking into account that in the strong-field ionization three-step model [1] there is a mapping between the ionization time and the momentum, and substituting $\langle\psi_{\mathbf{p}}^{(f)}(t)|\hat{\mathbf{p}}|\Psi(t)\rangle\rightarrow\mathbf{v}(t,t^{\prime})$, as well as $\langle 1|\hat{\mathbf{A}}(t)|0\rangle\rightarrow\langle\hat{\mathbf{A}}(t)\rangle\equiv-1/(2\pi\sqrt{\omega})\hat{\mathbf{e}}e^{i\omega t}$, we arrive at the simpleman Drude model representation for the spontaneous radiation of Eq. (22):
$\frac{dw_{\rm{D}}}{d\omega}=\frac{4\pi\omega^{2}}{c^{3}}\int dt^{\prime}\left|\int^{t_{f}+i\gamma/\omega}_{t^{\prime}+i\gamma/\omega}dt\mathbf{v}(t,t^{\prime})\cdot\langle\hat{\mathbf{A}}(t)\rangle\right|^{2}W(t^{\prime}).$
The results via the Drude model (via the current $J_{D}$) for the spontaneous radiation spectra are also presented in Fig. 3. The Drude model is qualitatively in accordance with $m_{1}$ in the low frequency region and confirms the enhancement of spectra in this frequency region in contrast to the behavior of the spectra based on the direct ionization $m_{0}$ amplitude. At large frequencies $\omega>6$ eV, when the radiation is dominated by interference of $m_{0}$ and $m_{1}$, the Drude model still surprisingly agrees quite well with the SFA-calculation.
III.3 Correlated electron and photon measurement
@@
The under-the-barrier recollisions are described by the $m_{2}$ amplitude. Our calculations show that its contribution to the coherent and spontaneous radiation spectral distribution is negligible, i.e., the radiation spectra do not carry information about under-the-barrier recollisions. Is there any other way to see this signature in the radiation properties? To answer this question, we discuss the correlated electron and photon measurement.
The spontaneous emission spectra in Fig. 3 have been calculated integrating over the final electron momentum. However, when the correlated electron and photon measurement is carried out in an experiment, one can access the electron momentum resolved photon spectra. The electron momentum resolved total photon yield in a half-cycle laser pulse in the case of the correlated measurement is defined as follows:
$\frac{dw}{d^{3}\mathbf{p}}=\frac{4\pi}{c^{3}}\int\omega^{2}d\omega\left|m_{0}(\mathbf{k},\mathbf{p})+m_{1}(\mathbf{k},\mathbf{p})\right|^{2}.$ (31)
The zeroth-order amplitude
$m_{0}(\mathbf{k},\mathbf{p})=\int dt\hat{\mathbf{A}}(t)\cdot\mathbf{m}_{p}(\mathbf{p},t),$ (32)
is calculated via the saddle-point integration in time $t$, where the saddle points are found numerically in a half-cycle sinusoidal laser pulse. For the calculation of the first-order amplitude, the Drude model for spontaneous emission is applied:
$m_{1}(\mathbf{k},p_{x},p_{\perp}=0)=\int^{t_{f}+i\gamma/\omega}_{t^{\prime}+i\gamma/\omega}dt\mathbf{v}(t,t^{\prime})\cdot\langle\hat{\mathbf{A}}(t)\rangle\sqrt{W(t^{\prime},p_{\perp}=0)}$
with the time-to-momentum mapping $p_{x}=-A(t^{\prime})$, and $W(t,p_{\perp}=0)=W(t)\kappa/(\pi|\mathbf{E}(t)|)$. The $t$-integration in $m_{1}$, and $\omega$-integration in Eq. (31) are carried out numerically.
We include the second-order correction to the emission amplitude $m_{2}$ heuristically, taking into account the correction in the ionization rate, where we replace the binding energy of the initial state $\kappa^{2}/2$ by $(\kappa^{2}/2)\{1+i\exp[-(2/3)(\kappa^{3}/|\mathbf{E}(t)|)]\}$ [50].
The correlated photon yield is presented in Fig. 4. The higher order corrections to the spectrum via $m_{0}$ and $m_{1}$ induce a shift of the position of the maximum. This is a result of the continuum motion. While in the given half-cycle laser pulse the photoelectron momentum peak is shifted from the zero momentum, the momentum shift $\delta p$ in the photoelectron momentum resolved photon yield is smaller, because it is additionally multiplied in Eq. (III.3) by the Fourier component of the ionizing electron velocity. The momentum shift approximately scales as $\delta p\sim(E_{0}/\kappa^{3})^{2/3}(E_{0}/\omega_{0})$ according to the fit to the numerical calculations.
The small contribution of the under-the-barrier recollisions via $m_{2}$ can be distinguished near the peak. However, this does not lead to any qualitative change of the spectrum.
Figure 5: The radiation spectrum at strong field ionization near the low-frequency region for the approximate parameters of the experiment of Ref. [15]: $E_{0}=0.1$, ${\cal A}=0.2$.
IV Comparison with experimental results
[…]
V Estimation of radiation intensities
Finally, let us give an order of magnitude estimations of the coherent and incoherent Brunel radiation. We proceed with the dipole radiation power [45]: